1. Find the current density and drift velocity of free electrons when the particles travel on a copper conductor with a square cross-section of 1.0 mm on the side and carries a current of 40 A to a 5 hp electric meter and its density of 6.0 x 1028 electrons/m3




2. Suppose the resistance is 1.47 Ω at a temperature of 20°C. Find the resistance at 0 degree Celsius.

1
Expert's answer
2022-03-09T15:55:11-0500

Solution:

"1. a=1\sdot10^{-3}m;"

"I=40A;"

"n=6.0\sdot10^{28}m^{-3};"

"j=\dfrac{I}{A};"

"A=a^2=1\sdot10^{-6}m^2;"

"j=\dfrac{40}{1\sdot10^{-6}}=40\sdot 10^6A/m^2;"

"v_d=\dfrac{I}{neA};"

"v_d=\dfrac{40}{6.0\sdot10^{28}\sdot1.6\sdot10^{-19}\sdot 1\sdot10^{-6}}=4.17\sdot10^{-3} m/s;"

"2. R=1.47\Omega;"

"T=20^oC;"

"T_0=0^oC;"

"\alpha=0.0043;"

"R_0-?;"

"R=R_0(1+\alpha\Delta T)" ;

"R_0=\dfrac{R}{1+\alpha \Delta T};"

"R_0=\dfrac{1.47}{1+0.0043\sdot20}=1.35\Omega;"


Answer:"1. j=40\sdot10^6 A/m^2; v_d=4.17\sdot10^{-3}m/s;"

"2. R_0=1.35\Omega."